OpenStudy (anonymous):

PLZ HELP! -2z-xy=x+7forx

OpenStudy (anonymous):

lets put all the \(x\) on the left, by adding \(xy\) to both sides you should get \[-2z=xy+x+7\] then subtract \(7\) from both sides, giving \[-2z-7=x+xy\] now at least we have all the \(x\)'s on one side

OpenStudy (anonymous):

i meant put them on the right, sorry

OpenStudy (anonymous):

now that we have all the \(x\) on the right, we can factor as \[-2z-7=(1+y)x\] and last step is to divide by \(1+y\) to get \[\frac{-2z-7}{1+y}=x\]

OpenStudy (anonymous):

so thats the answer

OpenStudy (anonymous):

that is what i get, yes. check my algebra, but i am pretty sure it is right

OpenStudy (anonymous):

can u help me on my other one